# Parsimonious Generative Machine Learning for Non-Gaussian Tail Modeling and Risk-Neutral Distribution Extraction

### 22 Feb 2024 v1

@@ Abstract
In financial modeling problems, non-Gaussian tails exist widely in many circumstances. Among them, the accurate estimation of risk-neutral distribution (RND) from option prices is of great importance for researchers and practitioners. A precise RND can provide valuable information regarding the market’s expectations, and can further help empirical asset pricing studies. This paper presents a parsimonious parametric approach to extract RNDs of underlying asset returns by using a generative machine learning model. The model incorporates the asymmetric heavy tails property of returns with a clever design. To calibrate the model, we design a Monte Carlo algorithm that has good capability with the assistance of modern machine learning computing tools. Numerically, the model fits Heston option prices well and captures the main shapes of implied volatility curves. Empirically, using S&P 500 index option prices, we demonstrate that the model outperforms some popular parametric density methods under mean absolute error. Furthermore, the skewness and kurtosis of RNDs extracted by our model are consistent with intuitive expectations.
Risk-Neutral Distribution,
Non-Gaussian Tails,
Generative Machine Learning,
No Arbitrage,
Calibration,
Implied Volatility Curves,
@@
[…]
, , and
@@ 1 Introduction
It has been extensively documented that non-Gaussian tails exist widely in financial markets (Rachev, 2003; Jondeau, Poon and Rockinger, 2007). Moreover, generally, the tails are asymmetric on the left and the right sides (Cont, 2001; McNeil, Frey and Embrechts, 2015). To avoid the concern of using Gaussian assumption, many advanced distributions were developed, such as the normal inverse Gaussian distribution (Barndorff-Nielsen, 1997), the hyperbolic distribution (Barndorff-Nielsen, 1977), the variance gamma distribution (Madan and Seneta, 1990), the generalized skewed $t$ distribution (Theodossiou, 1998), and so on. Most of them are defined via a relatively complicated density function, which restricts the flexible usage in numerical applications. While the advantage is the analytical tractability, the disadvantages are still significant, such as the inconvenient sampling, the difficult computation of the expectation of a function of the random variable, the further optimization of it, and the optimization of the conditional densities.
@@
Besides, the direct extraction of the risk-neutral distribution (RND), which is the distribution of the future underlying asset price/return in the risk-neutral world, is of great importance for both researchers and market practitioners. Theoretically, option prices reflect information about the financial market, represent the market’s expectations on the future, and provide a path for investments and risk management. The temporal behavior of the second moment of the market index’s RND is a barometer for monitoring the systemic risk or is the proxy of a risk factor. More specifically, risk-neutral higher moments or other distributional characteristics such as skewness, kurtosis, and quantiles derived from the estimated RND can have informational content or predictive power with economic significance in empirical asset pricing studies (Bali and Murray, 2013; Völkert, 2015; Stilger, Kostakis and Poon, 2017; Wong and Heaney, 2017; Chordia, Lin and Xiang, 2021; Fuertes, Liu and Tang, 2022). Furthermore, when RND is extracted from the market quotes of several options, we can use it to price other options for trading and hedging purposes. Given an accurate RND, any path-independent option can be priced precisely.
@@
Many researchers proposed various methods for recovering RND from market quotes of options. Christoffersen, Jacobs and Chang (2013) provided a comprehensive review of the literature and summarized the comparisons among these articles. Figlewski (2018) reviewed the major ideas in this field. Generally, the methods extracting RND include parametric methods, semi-parametric methods, and non-parametric ones, based on how they express the distribution.
@@
Parametric methods are the most popular ones, for there are lots of existing parametric distributions available for use. In Fabozzi, Tunaru and Albota (2009), researchers summarized parametric distributions including Weibull, generalized beta, and generalized gamma for recovering RND. Among the parametric methods, mixture models are the most widely used ones for their flexibility and parsimony, e.g., the double log-normal distribution in Bliss and Panigirtzoglou (2002). Semi-parametric methods include expansion ones such as Edgeworth expansion in Jarrow and Rudd (1982). Also, Madan and Milne (1994) used a Hermite polynomial approximation. By combining the Fourier cosine series method and Dirac Delta function with the Carr-Madan spanning formula, Cui and Yu (2021) and Cui and Xu (2022) presented their model-free methods to extract the RND. Albert, Herold and Muck (2023) extrapolated implied volatilities by parametrizing only the tail with Generalized Extreme Value (GEV) distribution.
@@
It is commonly known that RND can be derived from the second partial derivative of the call value with respect to the strike price. Therefore, a way to calculate an RND non-parametrically is smoothing and filling in the option price curve or implied volatility curve as a function of the strike price first, and then calculating the second partial derivative. The non-parametric methods also contain the use of neural networks for option pricing. Ackerer, Tagasovska and Vatter (2020) presented a neural network approach to fit the implied volatility surface. In contrast, many others concentrate on modeling the option price directly using a neural network with option parameters as inputs, as concluded in Ruf and Wang (2020). Besides, the spline method is also a non-parametric way, such as in Figlewski (2008); Reinke (2020). Figlewski (2008) proposed using a fourth-order spline to interpolate the observed implied volatilities. Orosi (2015) parametrized the call value as a function of the strike price, and the RND is derived using the second partial derivative again non-parametrically.
@@ 2.1 Some Existing Methods for RND Extraction
Generalized Beta Density (GB). Bookstaber and McDonald (1987) proposed the generalized beta distribution to model the risk-neutral distribution. Denote that $a>0$ is the power parameter, $b>0$ is the scale parameter, $v>0$ is the first beta parameter, and $w>0$ is the second beta parameter. Let $B$ be a beta random variable with parameters $v$ and $w$, then $X=b(B/(1-B))^{1/a}$ is a generalized beta with parameters $(a,b,v,w)$. For the complete density function which is much complicated, please refer to Bookstaber and McDonald (1987); Jondeau, Poon and Rockinger (2007).
Edgeworth Exapnsion (EW). As mentioned in Jondeau, Poon and Rockinger (2007), Jarrow and Rudd (1982) captured deviations from log-normality by an Edgeworth expansion of the risk-neutral distribution around the log-normal distribution. Letting $Q$ be the cdf of the price random variable $S_{T}$, Jarrow and Rudd (1982) showed that an Edgeworth expansion in densities for the true probability distribution $Q$ around the log-normal cdf $L$ can be written up to order 4 with cumulants. With some assumptions, the price of a European call option with strike price $K$ can be approximated. The implied volatility, skewness, and kurtosis can be estimated with nonlinear least squares using this expression.
@@ 3.1 An Existing Parsimonious GML Model
First proposed by Wu and Yan (2019) and further modified by Yan, Wu and Zhang (2019), researchers proposed a novel parametric quantile function to represent a univariate distribution for asset returns in the physical world. It is a monotonically increasing nonlinear transformation of the standard Gaussian quantile function $\Phi^{-1}(\alpha),~{}\alpha\in(0,1)$:
@@ 3.2.3 Tail Heaviness
Notice that plenty of distributions satisfy either of these two conditions above, and thus this proposition says that for many choices of $F$, any transform $f(x)$ satisfying $g_{1}(x)\to+\infty$ will yield a new distribution with a heavier tail. Furthermore, we state that by choosing appropriate $g_{1}(x)$ and $g_{2}(x)$, our proposed transformation can generate a new distribution $F_{2}$ whose tail is almost arbitrarily heavier than that of the base distribution $F_{1}$.
@@ 6 Conclusions
Appendix

### 04 Mar 2024 v2

@@ Abstract
In financial modeling problems, non-Gaussian tails exist widely in many circumstances. Among them, the accurate estimation of risk-neutral distribution (RND) from option prices is of great importance for researchers and practitioners. A precise RND can provide valuable information regarding the market’s expectations, and can further help empirical asset pricing studies. This paper presents a parsimonious parametric approach to extract RNDs of underlying asset returns by using a generative machine learning model. The model incorporates the asymmetric heavy tails property of returns with a clever design. To calibrate the model, we design a Monte Carlo algorithm that has good capability with the assistance of modern machine learning computing tools. Numerically, the model fits Heston option prices well and captures the main shapes of implied volatility curves. Empirically, using S&P 500 index option prices, we demonstrate that the model outperforms some popular parametric density methods under mean absolute error. Furthermore, the skewness and kurtosis of RNDs extracted by our model are consistent with intuitive expectations. More generally, the proposed methodology is widely applicable in data fitting and probabilistic forecasting.
@@
Risk-Neutral Distribution, Non-Gaussian Tails, Generative Machine Learning, No Arbitrage, Calibration, Implied Volatility Curves
journal: Journal of Econometrics
@@ 1 Introduction
It has been extensively documented that non-Gaussian tails exist widely in financial markets (Rachev, 2003; Jondeau et al., 2007). Moreover, generally, the tails are asymmetric on the left and the right sides (Cont, 2001; McNeil et al., 2015). To avoid the concern of using Gaussian assumption, many advanced distributions were developed, such as the normal inverse Gaussian distribution (Barndorff-Nielsen, 1997), the hyperbolic distribution (Barndorff-Nielsen, 1977), the variance gamma distribution (Madan and Seneta, 1990), the generalized skewed $t$ distribution (Theodossiou, 1998), and so on. Most of them are defined via a relatively complicated density function, which restricts the flexible usage in numerical applications. While the advantage is the analytical tractability, the disadvantages are still significant, such as the inconvenient sampling, the difficult computation of the expectation of a function of the random variable, the further optimization of it, and the optimization of the conditional densities.
@@
Besides, the direct extraction of the risk-neutral distribution (RND), which is the distribution of the future underlying asset price/return in the risk-neutral world, is of great importance for both researchers and market practitioners. Theoretically, option prices reflect information about the financial market, represent the market’s expectations on the future, and provide a path for investments and risk management. The temporal behavior of the second moment of the market index’s RND is a barometer for monitoring the systemic risk or is the proxy of a risk factor. More specifically, risk-neutral higher moments or other distributional characteristics such as skewness, kurtosis, and quantiles derived from the estimated RND can have informational content or predictive power with economic significance in empirical asset pricing studies (Bali and Murray, 2013; Völkert, 2015; Stilger et al., 2017; Wong and Heaney, 2017; Chordia et al., 2021; Fuertes et al., 2022). Furthermore, when RND is extracted from the market quotes of several options, we can use it to price other options for trading and hedging purposes. Given an accurate RND, any path-independent option can be priced precisely.
@@
Many researchers proposed various methods for recovering RND from market quotes of options. Christoffersen et al. (2013) provided a comprehensive review of the literature and summarized the comparisons among these articles. Figlewski (2018) reviewed the major ideas in this field. Generally, the methods extracting RND include parametric methods, semi-parametric methods, and non-parametric ones, based on how they express the distribution.
@@
Parametric methods are the most popular ones, for there are lots of existing parametric distributions available for use. In Fabozzi et al. (2009), researchers summarized parametric distributions including Weibull, generalized beta, and generalized gamma for recovering RND. Among the parametric methods, mixture models are the most widely used ones for their flexibility and parsimony, e.g., the double log-normal distribution in Bliss and Panigirtzoglou (2002). Semi-parametric methods include expansion ones such as Edgeworth expansion in Jarrow and Rudd (1982). Also, Madan and Milne (1994) used a Hermite polynomial approximation. By combining the Fourier cosine series method and Dirac Delta function with the Carr-Madan spanning formula, Cui and Yu (2021) and Cui and Xu (2022) presented their model-free methods to extract the RND. Albert et al. (2023) extrapolated implied volatilities by parametrizing only the tail with Generalized Extreme Value (GEV) distribution.
@@
It is commonly known that RND can be derived from the second partial derivative of the call value with respect to the strike price. Therefore, a way to calculate an RND non-parametrically is smoothing and filling in the option price curve or implied volatility curve as a function of the strike price first, and then calculating the second partial derivative. The non-parametric methods also contain the use of neural networks for option pricing. Ackerer et al. (2020) presented a neural network approach to fit the implied volatility surface. In contrast, many others concentrate on modeling the option price directly using a neural network with option parameters as inputs, as concluded in Ruf and Wang (2020). Besides, the spline method is also a non-parametric way, such as in Figlewski (2008); Reinke (2020). Figlewski (2008) proposed using a fourth-order spline to interpolate the observed implied volatilities. Orosi (2015) parametrized the call value as a function of the strike price, and the RND is derived using the second partial derivative again non-parametrically.
@@ 2.1 Some Existing Methods for RND Extraction
Generalized Beta Density (GB). Bookstaber and McDonald (1987) proposed the generalized beta distribution to model the risk-neutral distribution. Denote that $a>0$ is the power parameter, $b>0$ is the scale parameter, $v>0$ is the first beta parameter, and $w>0$ is the second beta parameter. Let $B$ be a beta random variable with parameters $v$ and $w$, then $X=b(B/(1-B))^{1/a}$ is a generalized beta with parameters $(a,b,v,w)$. For the complete density function which is much complicated, please refer to Bookstaber and McDonald (1987); Jondeau et al. (2007).
Edgeworth Exapnsion (EW). As mentioned in Jondeau et al. (2007), Jarrow and Rudd (1982) captured deviations from log-normality by an Edgeworth expansion of the risk-neutral distribution around the log-normal distribution. Letting $Q$ be the cdf of the price random variable $S_{T}$, Jarrow and Rudd (1982) showed that an Edgeworth expansion in densities for the true probability distribution $Q$ around the log-normal cdf $L$ can be written up to order 4 with cumulants. With some assumptions, the price of a European call option with strike price $K$ can be approximated. The implied volatility, skewness, and kurtosis can be estimated with nonlinear least squares using this expression.
@@ 3.1 An Existing Parsimonious GML Model
First proposed by Wu and Yan (2019) and further modified by Yan et al. (2019), researchers proposed a novel parametric quantile function to represent a univariate distribution for asset returns in the physical world. It is a monotonically increasing nonlinear transformation of the standard Gaussian quantile function $\Phi^{-1}(\alpha),~{}\alpha\in(0,1)$:
@@ 3.2.3 Tail Heaviness
Remark 1.
The distribution obtained in Equation (2) or (3) has a heavier right tail than Gaussian, because the cumulative function of Gaussian satisfies Condition b) and $u^{x}/A\to+\infty$. The left-tail side will have a similar conclusion. On the other hand, for any $n\geq 1$, the $n$-th moment of this distribution exists because it is not difficult to verify the existence of $\mathbb{E}[Z^{n}(u^{Z}/A+v^{-Z}/A+1)^{n}]$ when $Z$ is the standard Gaussian. Hence, the tails of this distribution are not so heavy.
Notice that plenty of distributions satisfy either of the two conditions in Proposition 6, and thus this proposition says that for many choices of $F$, any transform $f(x)$ satisfying $g_{1}(x)\to+\infty$ will yield a new distribution with a heavier tail. Furthermore, we state that by choosing appropriate $g_{1}(x)$ and $g_{2}(x)$, our proposed transformation can generate a new distribution $F_{2}$ whose tail is almost arbitrarily heavier than that of the base distribution $F_{1}$.
@@ 4.2 Monte Carlo Pricing with No-Aribtrage Condition
4.2.1 Cross-Maturity Arbitrage
An option’s value will decline over time. Or equivalently, the longer the time-to-maturity $\tau$, the greater the option value. Mathematically, the partial derivative of the option value with respect to $\tau$ should be positive. Or, there will be arbitrage opportunity existing. This requires that the RNDs for different $\tau$ satisfy some constraint.
To address it, we may need to calibrate the RNDs for different $\tau$ together, for example, 1 month, 3 months, 6 months, 1 year, and so on. Given the fact that the option prices quoted in the market are monotonically increasing with respect to $\tau$ almost all the time, we thereby take a simpler approach of post-processing here. Firstly, we separately calibrate the RNDs for different $\tau$. Then, we examine if the cross-maturity arbitrage is absent, or, if the option values priced with our RNDs are monotonically increasing with those $\tau$. If not, we tune or adjust the RND parameters $\sigma_{\tau},u_{\tau},v_{\tau}$ ($\mu_{\tau}$ is determined by them) with a slight amount, to eliminate the cross-maturity arbitrage. More specifically, we increase $\sigma_{\tau_{2}}$ slightly when the extracted RND at $\tau_{2}$ yields smaller option values than the RND at $\tau_{1}$ does, when $\tau_{2}>\tau_{1}$.
Indeed, in our numerical studies, separately calibrating the RNDs for different $\tau$ does not result in the violation of the cross-maturity arbitrage constraint. In other words, the post-processing does not happen even once in our experimental studies. However, the interpolation of RNDs for all $\tau\in[0,\mathcal{T}]$ will be a difficult task, and is out of the scope of this paper.
@@ 6 Conclusions
More generally, the proposed methodology is widely applicable in data fitting and probabilistic forecasting. It is more practical than many existing methods. The parameter estimation or learning given the dataset can be done via maximum likelihood or quantile regression.
Appendix A Some Proofs